Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Sun Feb 11 01:57:56 MET 2001) Success
Everything OK!

Effelsberg (A. Kraus at Mon Feb 12 09:44:13 MET 2001) Minor failures
Due to strong wind we had to stop the antenna between UT 0047-0118.

Torun (K.Borkowski at Mon Feb 12 09:50:44 MET 2001) Success
No known problems.

Onsala (John Conway at Mon Feb 12 10:04:16 MET 2001) Success
No known problems in w414a.

Westerbork (Tony Foley at Mon Feb 12 12:10:01 MET 2001) Minor failures
Known problem from one of the LOs going out of lock from 20:22 to 20:27 as a GSM phone came too close to the LO. The operator will rename nameless.

Jodrell_Bank (Alastair Gunn at Mon Feb 12 13:20:16 MET 2001) Minor failures
Many problems with telescope control software. Difficult to assess possible loss of data. Data should be mostly good.

Noto (C.Trigilio at Thu Feb 15 18:35:51 MET 2001) Success
No known problems.